Tom built a fence around his rectangular shaped pool. The posts are 2 yards apart, there are 4 posts along the width and 8 posts along the length. What
are the dimensions of the fence?
1
Expert's answer
2020-02-13T09:42:06-0500

/the dimensions of the fence is the perimeter fence

Prectangular="\Sigma" widths&lengths   (sum of side)

2*(widths+lengths)

from the first to the n-th posts (n-1) interval.

therefore the length in yards of width and side is respectively (4-1) * 2 = 6; (8-1) * 2 = 14; The perimeter is (6 + 14) * 2 = 40 yards